# Policy Prototyping for LLMs: Pluralistic Alignment via Interactive and Collaborative Policymaking

### v1

@@ Abstract
Emerging efforts in AI alignment seek to broaden participation in shaping model behavior by eliciting and integrating collective input into a policy for model finetuning. While pluralistic, these processes are often linear and do not allow participating stakeholders to confirm whether potential outcomes of their contributions are indeed consistent with their intentions. Design prototyping has long advocated for rapid iteration using tight feedback loops of ideation, experimentation, and evaluation to mitigate these issues. We thus propose policy prototyping for LLMs, a new process that draws inspiration from prototyping practices to enable stakeholders to collaboratively and interactively draft LLM policies. Through learnings from a real-world LLM policymaking initiative at an industrial AI lab, we motivate our approach and characterize policy prototyping with four guiding principles. Because policy prototyping emphasizes a contrasting set of priorities compared to previous approaches, we envision our approach to be a valuable addition to the methodological repertoire for pluralistic alignment.
@@ 1 Introduction and Background
Policies govern the world around us, from diplomatic relations on the world stage to resource allocation decisions in times of crisis. With the proliferation of products and services powered by large language models (LLMs), it is unsurprising that recent AI alignment and safety efforts have been increasingly invested in LLM policymaking [6, 9, 12, 16]. We refer to an LLM policy as text-based content describing acceptable and unacceptable model behaviors, along with any relevant context and definitions, that is then used to finetune and/or directly instruct the model [6, 31] [^1] In this paper, a policy refers to the set of rules, guidelines, and constraints that govern a model’s behavior and outputs. Note that the policies around which our work is centered are distinct from reinforcement learning policies, although both have similar goals in steering the model towards more desirable behaviors..
@@
Policy testing and evaluation with stakeholders is a well-known challenge identified long before the advent of LLMs [14, 17, 18, 20]. To tackle this, practitioners previously drew inspiration from design prototyping to propose policy prototyping [21, 23, 24], where policymakers iteratively collect and incorporate feedback on policies before publishing them. Specifically, developing a rough sketch of a policy (similar to a low fidelity prototype) to start allows stakeholders to identify desiderata and flaws earlier in the process [5, 30], while divergent-convergent thinking promoted by the double-diamond design process enables exploration of more alternatives before finalizing a policy [3]. Participatory design also provides methodological frameworks for empowering stakeholder engagement [8, 19, 26]. However, while conceptually appealing, policy prototyping is inherently difficult to achieve with traditional policies due to extended evaluation timescales and long feedback loops [7, 14]. LLM policymaking, however, is uniquely well-suited for policy prototyping, as LLM behaviors can be adapted quickly within tight feedback loops for proof-of-concept experimentation and evaluation (e.g., via system prompting [1]).
@@
In democratic societies, policies are the fruit of pluralistic input from diverse stakeholders. In a similar spirit, Huang et al., [16] incorporate public input into LLM policymaking by allowing a representative sample of the U.S. public to vote on and create statements in a “constitution” used to finetune a collectively-aligned LLM [2]. Feng et al. [10] propose collecting public input on “cases” that clarify ambiguous or vague statements in a constitution to improve the granularity of collective feedback. Despite the participatory intentions of these methods, it is not guaranteed that pluralistic stakeholder feedback is truly incorporated into model behavior as intended. This is due to the isolation of public input elicitation from model behavior adaptation in the alignment pipeline—stakeholders provide policy input, typically early on in the pipeline [2, 16]—without a means of directly experimenting with models that incorporate their input. Additionally, stakeholders may not see the impact of their input until the model is finetuned, tested, and released. This “participatory ceiling” [29] prevents stakeholders from verifying and iterating on their input to close the loop on their contributions.
@@
In this work, to address these limitations, we introduce policy prototyping for LLMs (henceforth “policy prototyping” for brevity), a new process for pluralistic alignment by which stakeholder groups can interactively and collaboratively prototype LLM policies, test resulting model behaviors, and resolve disagreements in real-time before a finalized policy is used for finetuning. We motivate and define guiding principles for policy prototyping with findings from a real-world LLM policymaking initiative in an industrial AI lab. Our guiding principles are meant to characterize this practice while still providing substantial flexibility for the process to be adapted when needed. We then discuss practical considerations of adopting policy prototyping—namely cost, scale, and tooling. Our work contributes an interdisciplinary avenue, bridging practices from policymaking and design, to enrich and complement existing approaches in pluralistic alignment.
@@ 2.2 Guiding Principles and Findings
GP2: Support synchronous collaboration and discussion. Unlike prior work that collected human feedback via asynchronous annotations [2, 13, 16, 22], our workshops had experts meet, discuss, and collaborate on policymaking synchronously. Experts unanimously found real-time collaboration to be enjoyable and productive. In E1’s words, “I found it hugely rewarding and beneficial personally and professionally […] I think we can get stuck in our heads because we’re working on our own so much.” E6 emphasized the support and learning opportunities afforded by collaboration: “[it was] very supportive having other voices in the back of your head […] it’s been incredible learning with everyone.” E9 found collaboration invaluable for surfacing new perspectives and broadening coverage of a broad domain: “there were times where I’m adamant that this is this, but someone else said something that just never occurred to me. And I think that’s why you need a *group* of experts.” The group managed to efficiently resolve some disagreements through real-time discussion, such as aligning on an interpretation of user intent behind a particular user query to an LLM or class of queries. These resolutions may take much longer to reach through asynchronous workflows. While this may not be true for all disagreements, it is clear that synchronous collaboration—currently underexplored in pluralistic alignment [28, 27]—has the potential to enhance both policy outcomes and experts’ policymaking experiences.
@@
GP4: Use scenarios as guiding artifacts. Scenario-based prototyping is a long-standing design practice [4, 15]. We found that experts engaged in critical discussions that led to nuanced policy considerations and suggestions when they deliberated with scenarios—example user queries within a specific domain [^3] Scenarios have also been referred to as “cases” in prior work [10].. Sample scenarios can be found in Appendix A.2. For E1, looking at scenarios helped them raise two key dimensions the model should consider in its response: “We need to ask clarifying questions, in particular to clarify the severity and the nature of the [user query]. Another dimension is to identify how long they’ve been [experiencing this].” E2 agreed with the need for a severity assessment, suggesting to present “an [urgency] rating scale on the scale of zero to 10” to the user as a simple first step. Adding on, E3 suggested eliciting the user’s financial ability to pay for the domain-specific service and making referrals accordingly: “There might be questions instead like, what is your financial ability to pay for [this service] right now? And if it’s within certain ranges, then you might make a community referral, like here’s some people in your area.” Scenario exploration also helped surface patterns in model responses and, through collective sensemaking, experts can then translate to behavioral rules for the model, as E7 describes: “I keep seeing this thing over and over and it’s incorrect, so that needs to be a rule.” In general, scenarios productively guided experts’ discussions and expanded opportunities for them to draw upon their expertise.
@@ 3 Discussion and Conclusion
Conclusion. We propose policy prototyping for LLMs as a new approach for pluralistic alignment. Policy prototyping draws from core ideas in design prototyping to empower stakeholder groups to collaboratively, interactively, and rapidly draft and test LLM policies in real time. Our approach grounds alignment in concrete, real-world experiences and human expertise elicited from rich, synchronous deliberation; its highly iterative nature contrasts with existing linear LLM policymaking processes that are designed to yield a high-fidelity, “production ready” policy directly. We envision policy prototyping to be used alongside existing pluralistic alignment methods—particularly earlier on in the alignment pipeline before a final policy is used to finetune a model—to maximize target alignment outcomes in a manner that also meaningfully incorporates human expertise and experiences.

### v2

@@ Abstract
Emerging efforts in AI alignment seek to broaden participation in shaping model behavior by eliciting and integrating collective input into a policy for model finetuning. While pluralistic, these processes are often linear and do not allow participating stakeholders to confirm whether potential outcomes of their contributions are indeed consistent with their intentions. Design prototyping has long advocated for rapid iteration using tight feedback loops of ideation, experimentation, and evaluation to mitigate these issues. We thus propose policy prototyping for LLMs, a new process that draws inspiration from prototyping practices to enable stakeholders to collaboratively and interactively draft LLM policies. Through learnings from a real-world LLM policymaking initiative at an industrial AI lab, we motivate our approach and characterize policy prototyping with four guiding principles. Because policy prototyping emphasizes a contrasting set of priorities compared to previous approaches, we envision our approach to be a valuable addition to the methodological repertoire for collaborative, pluralistic alignment.
@@ 1 Introduction and Background
Policies govern the world around us, from diplomatic relations on the world stage to resource allocation decisions in times of crisis. With the proliferation of products and services powered by large language models (LLMs), it is unsurprising that recent AI alignment and safety efforts have been increasingly invested in LLM policymaking [6; 9; 12; 16]. We refer to an LLM policy as text-based content describing acceptable and unacceptable model behaviors, along with any relevant context and definitions, that is then used to finetune and/or directly instruct the model [6; 31] [^1] In this paper, a policy refers to the set of rules, guidelines, and constraints that govern a model’s behavior and outputs. Note that the policies around which our work is centered are distinct from reinforcement learning policies, although both have similar goals in steering the model towards more desirable behaviors..
@@
Policy testing and evaluation with stakeholders is a well-known challenge identified long before the advent of LLMs [14; 17; 18; 20]. To tackle this, practitioners previously drew inspiration from design prototyping to propose policy prototyping [21; 23; 24], where policymakers iteratively collect and incorporate feedback on policies before publishing them. Specifically, developing a rough sketch of a policy (similar to a low fidelity prototype) to start allows stakeholders to identify desiderata and flaws earlier in the process [5; 30], while divergent-convergent thinking promoted by the double-diamond design process enables exploration of more alternatives before finalizing a policy [3]. Participatory design also provides methodological frameworks for empowering stakeholder engagement [8; 19; 26]. However, while conceptually appealing, policy prototyping is inherently difficult to achieve with traditional policies due to extended evaluation timescales and long feedback loops [7; 14]. LLM policymaking, however, is uniquely well-suited for policy prototyping, as LLM behaviors can be adapted quickly within tight feedback loops for proof-of-concept experimentation and evaluation (e.g., via system prompting [1]).
@@
In democratic societies, policies are the fruit of pluralistic input from diverse stakeholders. In a similar spirit, Huang et al., [16] incorporate public input into LLM policymaking by allowing a representative sample of the U.S. public to vote on and create statements in a “constitution” used to finetune a collectively-aligned LLM [2]. Feng et al. [10] propose collecting public input on “cases” that clarify ambiguous or vague statements in a constitution to improve the granularity of collective feedback. Despite the participatory intentions of these methods, it is not guaranteed that pluralistic stakeholder feedback is truly incorporated into model behavior as intended. This is due to the isolation of public input elicitation from model behavior adaptation in the alignment pipeline—stakeholders provide policy input, typically early on in the pipeline [2; 16]—without a means of directly experimenting with models that incorporate their input. Additionally, stakeholders may not see the impact of their input until the model is finetuned, tested, and released. This “participatory ceiling” [29] prevents stakeholders from verifying and iterating on their input to close the loop on their contributions.
@@
In this work, to address these limitations, we introduce policy prototyping for LLMs (henceforth “policy prototyping” for brevity), a new process for pluralistic alignment by which stakeholder groups can interactively and collaboratively prototype LLM policies, test resulting model behaviors, and resolve disagreements in real-time before a finalized policy is used for finetuning. We motivate and define guiding principles for policy prototyping with findings from a real-world LLM policymaking initiative in an industrial AI lab. Our guiding principles are meant to characterize this practice while still providing substantial flexibility for the process to be adapted when needed. We then discuss practical considerations of adopting policy prototyping—namely cost, scale, and tooling. Our work contributes an interdisciplinary avenue, bridging practices from policymaking and design, to enrich and complement existing approaches in collaborative, pluralistic alignment.
@@ 2.2 Guiding Principles and Findings
GP2: Support synchronous collaboration and discussion. Unlike prior work that collected human feedback via asynchronous annotations [2; 13; 16; 22], our workshops had experts meet, discuss, and collaborate on policymaking synchronously. Experts unanimously found real-time collaboration to be enjoyable and productive. In E1’s words, “I found it hugely rewarding and beneficial personally and professionally […] I think we can get stuck in our heads because we’re working on our own so much.” E6 emphasized the support and learning opportunities afforded by collaboration: “[it was] very supportive having other voices in the back of your head […] it’s been incredible learning with everyone.” E9 found collaboration invaluable for surfacing new perspectives and broadening coverage of a broad domain: “there were times where I’m adamant that this is this, but someone else said something that just never occurred to me. And I think that’s why you need a *group* of experts.” The group managed to efficiently resolve some disagreements through real-time discussion, such as aligning on an interpretation of user intent behind a particular user query to an LLM or class of queries. These resolutions may take much longer to reach through asynchronous workflows. While this may not be true for all disagreements, it is clear that synchronous collaboration—currently underexplored in pluralistic alignment [28; 27]—has the potential to enhance both policy outcomes and experts’ policymaking experiences.
@@
GP4: Use scenarios as guiding artifacts. Scenario-based prototyping is a long-standing design practice [4; 15]. We found that experts engaged in critical discussions that led to nuanced policy considerations and suggestions when they deliberated with scenarios—example user queries within a specific domain [^3] Scenarios have also been referred to as “cases” in prior work [10].. Sample scenarios can be found in Appendix A.2. For E1, looking at scenarios helped them raise two key dimensions the model should consider in its response: “We need to ask clarifying questions, in particular to clarify the severity and the nature of the [user query]. Another dimension is to identify how long they’ve been [experiencing this].” E2 agreed with the need for a severity assessment, suggesting to present “an [urgency] rating scale on the scale of zero to 10” to the user as a simple first step. Adding on, E3 suggested eliciting the user’s financial ability to pay for the domain-specific service and making referrals accordingly: “There might be questions instead like, what is your financial ability to pay for [this service] right now? And if it’s within certain ranges, then you might make a community referral, like here’s some people in your area.” Scenario exploration also helped surface patterns in model responses and, through collective sensemaking, experts can then translate to behavioral rules for the model, as E7 describes: “I keep seeing this thing over and over and it’s incorrect, so that needs to be a rule.” In general, scenarios productively guided experts’ discussions and expanded opportunities for them to draw upon their expertise.
@@ 3 Discussion and Conclusion
Conclusion. We propose policy prototyping for LLMs as a new approach for pluralistic alignment. Policy prototyping draws from core ideas in design prototyping to empower stakeholder groups to collaboratively, interactively, and rapidly draft and test LLM policies in real time. Our approach grounds alignment in concrete, real-world experiences and human expertise elicited from rich, synchronous deliberation; its highly iterative nature contrasts with existing linear LLM policymaking processes that are designed to yield a high-fidelity, “production ready” policy directly. We envision policy prototyping to be used alongside existing alignment methods—particularly earlier on in alignment pipelines before a final policy is used to finetune a model—to maximize target alignment outcomes in a manner that also meaningfully incorporates human expertise and experiences.